# Lost in Overlap: Exploring Watermark Collision in LLMs

### v1

@@ Abstract
The proliferation of large language models (LLMs) in generating content raises concerns about text copyright. Watermarking methods, particularly logit-based approaches, embed imperceptible identifiers into text to address these challenges. However, the widespread use of watermarking across diverse LLMs has led to an inevitable issue known as watermark collision during common tasks like question answering and paraphrasing. This study focuses on dual watermark collisions, where two watermarks are present simultaneously in the same text. The research demonstrates that watermark collision poses a threat to detection performance for detectors of both upstream and downstream watermark algorithms.
Lost in Overlap: Exploring Watermark Collision in LLMs
Yiyang Luo* Nanyang Technological University lawrenceluoyy@outlook.com                        Ke Lin*,† Tsinghua University leonard.keilin@gmail.com                        Chao Gu University of Science and Technology of China guch8017@mail.ustc.edu.cn
footnotetext: *Equal Contribution.footnotetext: Corresponding Author.
1 Introduction
As the text quality produced by large language models (LLM) advances, it tackles numerous practical challenges while raising many new issues. In particular, LLMs’ widespread generation of text on the Internet may increase rumors and raise concerns about text copyright Megías et al. (2021); Tang et al. (2023). Consequently, the identification and classification of machine-generated text have become critically significant. Watermarking methods can address this issue, making them increasingly prominent in current discussions and gathering growing attention.
Text watermarking involves embedding distinctive, imperceptible identifiers (watermarks) into written content. Nowadays, most methods are logit-based Kirchenbauer et al. (2023a); Liu et al. (2023); Zhao et al. (2023); Kuditipudi et al. (2023); Hu et al. (2023): they manipulate the output logits of LLMs during the generation process to achieve the goal of embedding watermarks. These watermarks are sophisticatedly designed to be robust yet discreet, ensuring content integrity and ownership preservation without compromising readability or meaning.
As LLMs flourish, the demand for watermarking rapidly escalates, leading to the application of various watermarking techniques across different LLMs. Consequently, particularly during common tasks such as question answering (QA) and paraphrasing, watermark collision becomes inevitable.
@@ 1 Introduction
The term watermark collision can be defined as instances of text with multiple watermarks occurring simultaneously (Fig. 1). While many methods Liu et al. (2023); Zhao et al. (2023); Kuditipudi et al. (2023) claim resilience against paraphrase attacks, none have been specifically tested for watermark collisions. As part of this study, we examine dual watermark collisions in which only two watermarks are involved in the same textual sample.
Figure 2: The method pipeline. $T_{raw}$, $T_{W}$, and $T_{P}$ denote the original input text data, text data with a single watermark, and text with a dual watermark, respectively. For a thorough comparison, we also introduce $T_{P}^{\prime}$, containing only a single watermark. $D_{W}$ and $D_{P}$ represent the watermark detectors for the watermarker $W$ and paraphraser $P$, respectively.
@@
The goal of this paper is to discuss watermark collisions and their impact on LLMs. Our contributions are as follows:
  • We propose a novel variant of paraphrase attacks that could easily remove existing watermarks from the text.
  • We find that the strength of the overlapping watermark impacts detection performance. Upstream and downstream watermarks are generally competing for detection accuracy, with one being stronger and the other weaker.
@@
2 Related Work
@@
Watermarks in the LLM era can be categorized into two categories Liu et al. (2024): post-processing and generation-based techniques. While post-processing techniques Abdelnabi and Fritz (2020); Yang et al. (2022) modify the text and preserve its semantics, generation-based methods incorporate watermarks into the text generation process to achieve better results. Kirchenbauer et al. (2023a) proposed a logit-based greenlist mechanism based on prior token hashes. Liu et al. (2023) introduced a watermark model to generate semantic-preserving logits during text generation. Zhao et al. (2023) simplified the Kirchenbauer et al.’s scheme by using a fixed Green-Red split and achieved greater robustness. Christ et al. (2023) and Kuditipudi et al. (2023) aims to design watermark techniques that are more robust and secure.
@@
Even though these methods have been designed to be more robust against such paraphrase attacks Kirchenbauer et al. (2023b), these paraphrase attacks often use unwatermarked LLMs, e.g. DIPPER Krishna et al. (2023) and GPT-3.5 Brown et al. (2020). Therefore, it is unclear what impact one watermark technique has over another, which is the motivation and objective of this study.
3 Method
We design a pipeline consisting of a watermarker and a paraphraser to simulate the collisions between different watermarks:
  • Watermarker $W$ creates watermarked texts $T_{W}$ by utilizing an LM to generate content based on specific corpus as context.
  • Paraphraser $P$ paraphrases the watermarked texts $T_{W}$ with different watermarks, i.e. generated by different methods or keys.
@@ 3 Method
As shown in Fig. 2, we first generate text data $T_{W}$ using Watermarker $W$. Then the text data will be paraphrased by paraphraser $P$ with different watermarks to generate paraphrased text data $T_{P}$. Paraphrased texts $T_{P}$ are intended to contain dual watermarks simultaneously, and the impact of watermark collisions is then investigated.
Furthermore, we also generate unwatermarked texts $T_{W}^{\prime}$ and $T_{P}^{\prime}$ as baselines using watermarkers and paraphrasers. Section 4.1 and Appendix A provide details regarding pipeline setup.
$W$ $P$ $\text{KGW}_{weak}$ $\text{SIR}_{weak}$ $\text{PRW}_{weak}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{weak}$ 52.80 19.80 3.40 90.09 41.10 48.00
$\text{SIR}_{weak}$ 55.74 25.10 / / 41.05 19.80
$\text{PRW}_{weak}$ 37.00 28.20 22.30 79.56 26.60 41.20
@@
$W$ $P$ $\text{KGW}_{strong}$ $\text{SIR}_{strong}$ $\text{PRW}_{strong}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{weak}$ 4.10 97.40 0.20 92.77 6.20 99.90
$\text{SIR}_{weak}$ 12.45 96.70 / / 13.22 97.50
$\text{PRW}_{weak}$ 14.60 96.70 29.40 91.91 9.30 99.90
@@
$W$ $P$ $\text{KGW}_{weak}$ $\text{SIR}_{weak}$ $\text{PRW}_{weak}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{strong}$ 68.20 21.50 9.00 79.26 56.00 44.40
$\text{SIR}_{strong}$ 61.55 21.00 / / 45.58 14.00
$\text{PRW}_{strong}$ 67.90 18.30 32.00 65.31 44.20 41.70
@@
$W$ $P$ $\text{KGW}_{strong}$ $\text{SIR}_{strong}$ $\text{PRW}_{strong}$
$D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
$\text{KGW}_{strong}$ 7.00 94.10 0.80 90.52 14.80 99.70
$\text{SIR}_{strong}$ 10.98 97.40 / / 12.58 97.10
$\text{PRW}_{strong}$ 14.60 96.70 37.10 79.00 11.20 99.30
@@
Table 1: TPR of the paraphrased text $T_{P}$ with dual watermarks when $\text{FPR}=1\%$. $W$ and $P$ represent the watermarker and paraphraser, respectively. $D_{W}$ and $D_{P}$ represent the detector of watermarker and paraphraser.
$W$ $T\&D$ ($T_{W}$, $D_{W}$) ($T_{P}^{\prime}$, $D_{W}$) ($T_{P}$, $D_{W}$)
$\text{KGW}_{weak}$ 99.90 71.65 52.80
$\text{KGW}_{strong}$ 100.00 81.95 68.20
$\text{SIR}_{weak}$ 87.90 59.05 55.74
$\text{SIR}_{strong}$ 94.30 67.60 61.55
$\text{PRW}_{weak}$ 95.40 49.25 37.00
$\text{PRW}_{strong}$ 99.70 75.05 67.90
Table 2: TPR of detecting watermark using $D_{W}$ with paraphrase attacks when $\text{FPR}=1\%$. ($T_{W}$, $D_{W}$) represents the TPR before paraphrasing attacks. Meanwhile, ($T_{P}^{\prime}$, $D_{W}$) denotes the TPR after paraphrasing attacks without watermarks, and ($T_{P}$, $D_{W}$) denotes the TPR after paraphrasing attacks with a watermark.
4 Experiments
4.1 Experiment Setup
@@
We utilize the C4 dataset Raffel et al. (2020) as the context for text generation with a maximum of 128 tokens. In the capacity of the watermarker, we employ both LLaMA-2-13B Touvron et al. (2023) and OPT-1.3B Zhang et al. (2022), while for paraphrasing, we exclusively utilize LLaMA-2-13B. Appendix A shows more details of the experiments.
@@
Our experiments were conducted using several previous watermark methods as baselines. The watermark strength of each method is tuned separately for weak and strong watermarks. Methods used include: (a) KGWfrom Kirchenbauer et al. (2023a), $d=2,5$ for weak and strong settings; (b) SIRfrom Liu et al. (2023), $d=2,5$ for weak and strong settings; (c) PRWfrom Zhao et al. (2023), $s=2,5$ for weak and strong settings.
@@ 4.2 Evaluation Metrics
Given the rapid development of the Watermarking field, there is currently no consensus on metrics, with different methods employing varied evaluation criteria. Moreover, many approaches utilize a detection threshold to categorize text as watermarked. Implementing all metrics for each method is impractical and biased. Following the approach of Zhao et al. (2023), we opt for a fair evaluation by avoiding the influence of threshold settings. Our final detection metrics include false positive rates (FPR) and true positive rates (TPR). We specifically set FPR values at 1%, 5%, and 10%, adjusting the detector’s thresholds accordingly. This ensures a consistent and unbiased assessment of watermarking methods. We only present $\text{FPR}=1\%$ here and refer to Appendix B for more results.
4.3 Experimental Results
@@
Paraphrase attacks with watermarks tend to be stronger than those without.
In Table 2, we present the detection accuracy for various baseline watermark algorithms with and without the occurrence of watermark collision ($T_{P}$ and $T_{P}^{\prime}$, respectively). A noteworthy decline in detection accuracy is observed when watermarks are introduced in the context of paraphrase attacks.
Different watermarks overlapping impair detection accuracy, with one being stronger and the other weaker.
@@
It is discernible that the watermark collision adversely impacts the integrity of both the watermarker and the paraphraser, leading to observable damage in their respective watermarks. According to the settings in Table 1, there is a strong competition between overlapping watermarks. As the watermarker attempts to maintain its detection accuracy, the paraphraser’s detection accuracy decreases. SIR-SIR is not listed since the SIR does not allow the user to choose a key to create a different distribution, and therefore the watermark collision will not occur.

### v2

# Lost in Overlap: Exploring Logit-based Watermark Collision in LLMs

@@ Abstract
The proliferation of large language models (LLMs) in generating content raises concerns about text copyright. Watermarking methods, particularly logit-based approaches, embed imperceptible identifiers into text to address these challenges. However, the widespread usage of watermarking across diverse LLMs has led to an inevitable issue known as watermark collision during common tasks, such as paraphrasing or translation. In this paper, we introduce watermark collision as a novel and general philosophy for watermark attacks, aimed at enhancing attack performance on top of any other attacking methods. We also provide a comprehensive demonstration that watermark collision poses a threat to all logit-based watermark algorithms, impacting not only specific attack scenarios but also downstream applications.
Introduction
As the quality of text produced by large language models (LLMs) advances, it addresses numerous practical challenges while raising many new issues. In particular, the widespread generation of text by LLMs on the Internet may increase the spread of rumors and raise concerns about text copyright (Megías et al. 2021; Tang et al. 2023). Consequently, the identification and classification of machine-generated text have become critically significant. Watermarking techniques for LLMs can help to tackle these problems, leading to their rising importance in ongoing conversations and attracting increasing interest globally.
Text watermarking involves embedding distinctive, imperceptible identifiers (watermarks) into written content. Nowadays, most methods are logit-based (Kirchenbauer et al. 2023a; Liu et al. 2023a; Zhao et al. 2023; Kuditipudi et al. 2023; Hu et al. 2023): they manipulate the output logits of LLMs during the generation process using distinct but consistently logit-based strategies to embed watermarks successfully. Utilizing the power of LLMs ensures that such adjustings for probabilistic distribution are seamlessly integrated into the generated content without compromising the overall quality or coherence of the text (Lin et al. 2024). These watermark methods are sophisticatedly designed to be robust yet discreet, ensuring content integrity and ownership preservation without compromising readability or meaning.
@@ Introduction
However, as watermarking techniques proliferate, watermark collision becomes inevitable with the increasing application of watermarks. The term watermark collision can be defined as instances where the text contains multiple watermarks simultaneously (Fig. 1). This is particularly inevitable during tasks that may require the collaboration of multiple LLMs, such as paraphrasing and translation. While many methods (Liu et al. 2023a; Zhao et al. 2023; Kuditipudi et al. 2023) claim resilience against paraphrase attacks, none have been specifically tested for watermark collisions. Hence, we examine the underlying mathematical principles, employ watermark collision as an attack strategy for all watermarking techniques that utilize logits, and evaluate its effectiveness in conjunction with multiple traditional attack methods enhanced through the incorporation of watermark collision.
@@
In summary, this paper proposes a new watermark attack philosophy for all logit-based watermarks in LLMs. Our contributions are as follows:
  • We propose a novel philosophy for watermark attacks that can effectively remove existing watermarks from text. This approach can be integrated with various traditional attack methods to enhance their performance.
  • We find that the strength of overlapping watermarks impacts detection performance. Upstream and downstream watermarks generally compete for detection accuracy, with one being stronger and the other weaker.
@@
Related Work
Text watermarking.
Modern text watermarking techniques can be classified into two categories: modification-based and generation-based watermarks (Liu et al. 2024). Modification-based watermarking, also known as watermarks for existing text, consists of altering an existing text to create a watermarked text. Most of the modification-based techniques can be classified as either lexical (Topkara, Topkara, and Atallah 2006b; Yang et al. 2022; Munyer and Zhong 2023; Sato et al. 2023) or syntactic methods (Atallah et al. 2001; Topkara, Topkara, and Atallah 2006a; Meral et al. 2009), based on rules, classical machine learning or deep neural models.
LLM watermarking.
While modification-based techniques (Abdelnabi and Fritz 2020; Yang et al. 2022) modify the text and preserve its semantics, generation-based methods apply watermarks into the text generation process to achieve better results, enabling smoother integration with LLMs. Watermark injection can be carried out during either the training phase or the inference phase.
During the training of LLMs, watermarks are inserted into the training data to intentionally alter the results of LLMs for certain inputs (Liu et al. 2023b; Sun et al. 2022). The main objective of training time watermarking is to protect dataset copyrights from unauthorized usage (Tang et al. 2023; Sun et al. 2023). Despite being able to embed watermarks in LLMs, training-time watermarking has significant limitations, including limited payload capacity, restricted trigger conditions, and significant training overhead.
@@
For inference-time watermarking, Kirchenbauer et al. (2023a) proposed a logit-based greenlist mechanism based on prior token hashes. Liu et al. (2023a) introduced a watermark model to generate semantic-preserving logits during text generation. Zhao et al. (2023) simplified the Kirchenbauer et al.’s scheme by using a fixed Green-Red split and achieved greater robustness. Christ, Gunn, and Zamir (2023); Kuditipudi et al. (2023); Fu et al. (2024) aim to design watermark techniques that are more robust and secure. Yoo, Ahn, and Kwak (2023); Boroujeny et al. (2024) introduce watermarking techniques with increased payload capacity for arbitrary binary data. Zhu et al. (2024) enhances the efficiency and quality of watermarking by embedding dual secret patterns.
@@
Even though these methods have been designed to be more robust against attacks such as paraphrase attacks (Kirchenbauer et al. 2023b), back-translation attacks (He et al. 2024) and mask-and-fill attacks (Lyu, Huang, and Kong 2023), these attacks often use unwatermarked LLMs, e.g. DIPPER (Krishna et al. 2023) and GPT-3.5 (Brown et al. 2020). Prior research in the pre-LLM era has mentioned potential risks associated with multiple watermarks (Tanha et al. 2012). Nevertheless, the effects of one watermarking technique on another in the context of LLM watermarking remain unclear, which is the motivation for this study.
Method
Figure 2: The collision pipeline. $T_{W}$ denotes text with the first watermark $W$, where $T_{C}$ denotes text with dual watermarks from a different collider $C\in\{P,R,M\}$. Unwatermarked text generated from $W$ and $C$ is denoted as $T_{W^{\prime}}$ and $T_{C^{\prime}}$. $T_{C}$ and $T_{C^{\prime}}$ are then examined by $D_{W}$ and $D_{C}$ to determine the presence of watermark $W$ and $C$. Texts in red and green are visualization samples of the red-green list showing the original watermark $W$.
Principle of Watermark Collision
The detection process for logit-based watermarking methods relies on the null hypothesis testing. A well-known example is the null hypothesis of KGW (Kirchenbauer et al. 2023a):
$H_{0}:{}$ The text sequence is generated with (1)
no knowledge of the red list rule.
Since the words of red list are chosen randomly, a natural writer is expected to sample words both from red and green list, whereas the watermarked model produces words only from green list. In practice, effective detection typically requires the text to contain sufficient words from a runtime-generated green list. This requirement ensures that the generated text adheres to a specific probability distribution. The detection algorithm checks whether the distribution of words in the text conforms to the green-list distribution.
Let $\mathbf{P}(T)$ represent the probability distribution of a certain text $T$. The detection process verifies if $\mathbf{P}(T)$ follows the null hypothesis $H_{0}$, which can be formulated as the follows ($\bm{\delta}(\cdot)$ is the dirichlet function):
$\bm{\delta}[T\text{ is watermarked}]=\begin{cases}0&\mathbf{P}(T)\approx\mathbf{P}(H_{0})\\ 1&\mathbf{P}(T)\not\approx\mathbf{P}(H_{0})\end{cases}$ (2)
From a probabilistic perspective, a watermark $w$ can be detected only when text $T$ follows a watermarked distribution $\mathbf{P}_{w}$. However, each watermarking method has an independent null hypothesis $H_{0}$ and therefore creates different word distributions in the generation process. When an additional watermark $w^{\prime}$ is applied, it imposes a new distribution $\mathbf{P}_{w^{\prime}}$ on the text. If more watermarks are added, a series of distributions are obtained: $(\mathbf{P}_{w^{(0)}},\mathbf{P}_{w^{(1)}},\mathbf{P}_{w^{(2)}},\dotsc,\mathbf{P}_{w^{(n)}})$, where $\mathbf{P}_{w^{(0)}}=\mathbf{P}_{w}$. Introducing each new watermark modifies the word probabilities in $T$, thus altering the overall distribution. Specifically, any distinct watermarks $w^{(i)}$ and $w^{(j)}$ should have different distributions $\mathbf{P}_{w^{(i)}}$ and $\mathbf{P}_{w^{(j)}}$ to ensure that they are not incorrectly detected by each other. As a result, the text $T$ no longer strictly follows the original distribution $\mathbf{P}(T)$, nor does it fully conform to any of the subsequent distributions $\mathbf{P}_{w^{(0)}},\mathbf{P}_{w^{(1)}},\mathbf{P}_{w^{(2)}},\dotsc,\mathbf{P}_{w^{(n)}}$. Therefore, the combination of multiple watermarks can be described as a transformation of the original watermark into an entangled one:
$\mathbf{P}_{\text{entangled}}=f(\mathbf{P}_{w^{(0)}},\mathbf{P}_{w^{(1)}},\mathbf{P}_{w^{(2)}},\dotsc,\mathbf{P}_{w^{(n)}})$ (3)
Here, $f$ represents the complex transformation function resulting from the sequential application of multiple watermarks. This new distribution $\mathbf{P}_{\text{entangled}}$ is not merely a simple combination but a new complex distribution that emerges from the (indirect) interaction of all applied watermarks.
Since the detection process relies on identifying patterns consistent with $\mathbf{P}_{w^{(i)}}$, the introduction of $\mathbf{P}_{\text{entangled}}$ causes the final text distribution to no longer conform to any of these patterns. The detection algorithm may not be able to detect the watermark in the text due to a shift in word distribution from $P_{w^{(i)}}$, leading to watermark collisions.
Pipeline Design
To prove the existence of watermark collisions, we design pipelines with three main components: watermarker, colliders, and detectors.
Watermarker
Watermarker $W$ generates watermarked texts $T_{W}$ by using a language model (LM) to create content based on a specific corpus as context. As illustrated in Fig. 2, we first produce the watermarked text data $T_{W}$ with Watermarker $W$. Additionally, we generate unwatermarked text $T_{W^{\prime}}$ using the same context and prompt as $T_{W}$ for further comparisons. Experiments Section and Appendix provide details regarding the watermarker setup.
Colliders
Colliders $C$ are designed to attack the watermark created by the watermarker using collision techniques. There are three distinct colliders that apply such collision attacks through traditional attack methods, namely paraphraser, back-translator, and mask-and-filler.
Paraphrase Collider.
@@
Paraphraser $P$ rephrases the watermarked texts $T_{W}$ with different watermarks, i.e. generated by different methods or keys, to generate paraphrased text data $T_{P}$, which are intended to contain dual watermarks simultaneously. Furthermore, we also generate texts $T_{P}^{\prime}$ using the same paraphraser but without a watermark, denoted as $P^{\prime}$, for further comparison.
Back-translation Collider.
Translator $R$ translates the watermarked texts $T_{W}$ to other languages and then translates back to their original language with watermarks. As shown in Fig. 2, we first paraphrase the original text data $T_{W}$ using Watermarker $W$. Then the text data will be translated and back-translated by Translator $R$ with different watermark settings to generate text data $T_{R}$ which are intended to contain dual watermarks simultaneously. Furthermore, we also generate texts $T_{R}^{\prime}$ using the same translator but without a watermark, denoted as $R^{\prime}$, for further comparison.
It is important to note that the KGW-based method is essentially ineffective against back-translation attacks due to the incapability of capturing the contextual semantics. Thus, they have been excluded from this pipeline. We choose French as the pivot language in the back-translation. Details are in Experiments Section.
Mask-and-fill Colliders.
Mask-and-filler (MnF) $M$ is specifically designed for mask-and-fill attacks. The MnF attack method is commonly used with masked language models, e.g., BERT-based models. For our study, we opted for RoBERTaLARGE as the base model. As shown in Fig. 2, we first generate watermarked text data $T_{W}$ using Watermarker $W$. Then the text data will be mask-and-filled by MnF $M$ with different watermark settings to generate text data $T_{M}$ which are intended to contain dual watermarks simultaneously. Additionally, we create texts $T_{M}^{\prime}$ by applying the same MnF but excluding the watermark, denoted as $M^{\prime}$. These will be used as baseline texts for comparison.
Detectors
As demonstrated in Fig.2, there are four detectors, each tailored to identify a specific type of watermark. Detector $D_{P}$ targets watermarks in paraphrasers, $D_{R}$ focuses on those in translators, and $D_{M}$ is for watermarks in the MnF process. Detector $D_{W}$ aims to identify the original watermark embedded by the watermarker. By comparing the results from these detectors, we can assess the effectiveness of the attacks with or without additional watermarks.
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
KGWwk 99.90 71.65 52.80 19.80 41.10 48.00 3.40 90.09
PRWwk 95.40 49.25 37.00 28.20 26.60 41.20 22.30 79.56
SIRwk 87.90 59.05 55.74 25.10 41.05 19.80 / /
@@
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
KGWwk 99.90 71.65 4.10 97.40 6.20 99.90 0.20 92.77
PRWwk 95.40 49.25 14.60 96.70 9.30 99.90 29.40 91.91
SIRwk 87.90 59.05 12.45 96.70 13.22 97.50 / /
@@
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
KGWsg 100.00 81.95 68.20 21.50 56.00 44.40 9.00 79.26
PRWsg 99.70 75.05 67.90 18.30 44.20 41.70 32.00 65.31
SIRsg 94.30 67.60 61.55 21.00 45.58 14.00 / /
@@
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$ $D_{W}$ $D_{P}$
KGWsg 100.00 81.95 7.00 94.10 14.80 99.70 0.80 90.52
PRWsg 99.70 75.05 14.60 96.70 11.20 99.30 37.10 79.00
SIRsg 94.30 67.60 10.98 97.40 12.58 97.10 / /
@@
Table 1: TPR of the paraphrased text $T_{P}$ with dual watermarks when $\text{FPR}=1\%$. $W$ and $P$ represent the watermarker and paraphraser, respectively. $D_{W}$ and $D_{P}$ represent the detector of the watermarker and paraphraser. $\varnothing$ indicates that no paraphrasing process is applied to the text, and its corresponding column represents the result of using $D_{W}$ to detect watermark $W$ in $T_{W}$. $P^{\prime}$ represents paraphrasing $T_{W}$ without watermark, as mentioned in Fig. 2.
Experiments
Experiment Setup
@@
We utilize the C4 dataset (Raffel et al. 2020) as the context for text generation with a maximum of 128 tokens. For watermarker and paraphraser, we employ LLaMA-2-13B (Touvron et al. 2023), Qwen2-7B (Bai et al. 2023) and OPT-1.3B (Zhang et al. 2022). For back-translator, we use LLaMA-2-13B and Qwen2-7B. For mask-and-filler, we exclusively utilize RoBERTaLARGE (Liu et al. 2019) as the masked language model. We only present results of LLaMA-2-13B here and refer to the Appendix for more details of the experiments.
@@
[…]
@@
Given the rapid development of the watermarking field, there is currently no consensus on metrics, with different methods employing varied evaluation criteria (Tu et al. 2023). Moreover, some approaches utilize a detection threshold to categorize text as watermarked, while others differ. Implementing all possible metrics for each method is impractical and biased.
Following the approach of Zhao et al. (2023), we opt for a fair evaluation by avoiding the influence of threshold settings. Our final detection metrics include false positive rates (FPR) and true positive rates (TPR). We specifically set FPR values at 1%, 5%, and 10%, adjusting the detector’s thresholds accordingly. This ensures a consistent and unbiased assessment of watermarking methods. We only present results when $\text{FPR}=1\%$ here and refer to Appendix for comprehensive results.
$W$ $R$varnothing $R^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$
KGWwk 99.90 44.80 41.90 9.20 34.20 13.70 31.40 5.15
PRWwk 95.40 32.90 34.10 7.50 21.80 12.70 25.20 5.73
SIRwk 87.90 5.60 4.75 7.70 6.12 11.50 / /
(a) weak $W$, weak $R$
$W$ $R$varnothing $R^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$
KGWwk 99.90 44.80 28.50 69.50 18.50 58.50 3.90 93.50
PRWwk 95.40 32.90 23.00 68.00 14.30 54.60 10.80 92.87
SIRwk 87.90 5.60 4.28 67.40 4.40 41.60 / /
(b) weak $W$, strong $R$
$W$ $R$varnothing $R^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$
KGWsg 100.00 69.30 67.80 8.10 61.90 13.60 55.30 4.09
PRWsg 99.70 63.90 64.80 9.60 55.20 10.20 50.60 3.56
SIRsg 94.30 3.00 1.74 9.40 3.82 6.70 / /
(c) strong $W$, weak $R$
$W$ $R$varnothing $R^{\prime}$ KGWwk PRWwk SIRwk
$D_{W}$ $D_{W}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$ $D_{W}$ $D_{R}$
KGWsg 100.00 69.30 50.00 68.90 38.80 53.00 7.50 91.70
PRWsg 99.70 63.90 50.10 67.90 36.60 42.40 20.10 90.44
SIRsg 94.30 3.00 4.15 74.40 2.51 22.50 / /
(d) strong $W$, strong $R$
Table 2: TPR of the back-translated text $T_{R}$ with dual watermarks when $\text{FPR}=1\%$. Similar to Table 1. Data annotated with underline indicate abnormal data points.
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
KGWwk 7.36 6.96 8.60 6.58 12.28
PRWwk 6.92 6.69 9.77 6.24 12.15
SIRwk 10.14 7.15 11.04 7.09 /
(a) weak $W$, weak $R$
$W$ $P$varnothing $P^{\prime}$ KGWsg PRWsg SIRsg
KGWwk 7.36 6.96 15.67 5.05 14.08
PRWwk 6.92 6.69 13.85 5.22 10.25
SIRwk 10.14 7.15 13.31 5.12 /
(b) weak $W$, strong $R$
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
KGWsg 13.18 7.44 14.16 8.39 13.55
PRWsg 8.31 7.61 10.53 6.54 12.00
SIRsg 12.30 7.72 12.75 7.80 /
(c) strong $W$, weak $R$
$W$ $P$varnothing $P^{\prime}$ KGWsg PRWsg SIRsg
KGWsg 13.18 7.44 12.80 5.60 12.03
PRWsg 8.31 7.61 11.15 4.91 9.22
SIRsg 12.30 7.72 11.29 4.99 /
(d) strong $W$, strong $R$
Table 3: PPL of the paraphrased $T_{P}$ with dual WMs.
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
KGWwk 7.36 8.82 9.00 8.29 10.16
PRWwk 6.92 8.21 8.52 8.02 9.39
SIRwk 10.14 9.82 10.37 9.14 /
(a) weak $W$, weak $R$
$W$ $P$varnothing $P^{\prime}$ KGWsg PRWsg SIRsg
KGWwk 7.36 8.82 11.86 9.63 33.35
PRWwk 6.92 8.21 11.15 8.84 30.57
SIRwk 10.14 9.82 13.39 9.92 /
(b) weak $W$, strong $R$
$W$ $P$varnothing $P^{\prime}$ KGWwk PRWwk SIRwk
KGWsg 13.18 12.74 12.88 11.41 14.91
PRWsg 8.31 9.02 9.29 8.33 11.11
SIRsg 12.30 10.05 10.91 8.72 /
(c) strong $W$, weak $R$
$W$ $P$varnothing $P^{\prime}$ KGWsg PRWsg SIRsg
KGWsg 13.18 12.74 17.58 11.83 36.84
PRWsg 8.31 9.02 12.59 9.60 27.98
SIRsg 12.30 10.05 14.78 9.18 /
(d) strong $W$, strong $R$
Table 4: PPL of the back-translated $T_{P}$ with dual WMs.
Experimental Results & Analysis
Watermark collision is compatible with the majority of existing attacks.
Tables 1, 2, and 5 demonstrate that watermark collision is feasible for all selected attacks, including paraphrase, back-translation, and mask-and-fill attacks. Watermark collision is commonly found in attacks involving auto-regressive text generation methods such as paraphrasing and back-translation. Mask-and-fill attacks are ineffective as they cannot completely change the distribution of words in a sentence.
Watermark collision will not degrade the text quality.
As shown in Tab. 3 and 4, we present perplexity before and after attacks, both with and without collisions. We use LLaMA-2-13B as the backbone for perplexity calculation, which is the same model for the initial generated text. As evidence, the text quality remains largely stable post-attack, and most collisions did not result in significant declines in text quality, indicating the potential value of collision as a methodology of attack.
@@
Watermark attacks with watermarks collision tend to be stronger than those without.
@@
In Table 1 and 2, we present the detection accuracy for various baseline watermark algorithms with and without the occurrence of watermark collision ($T_{P}$ and $T_{P}^{\prime}$, respectively). A noteworthy decline in detection accuracy is observed when watermarks are introduced in the context of traditional attacks such as paraphrase attacks and back-translation attacks. According to the settings in Table 1 and 2, there is a strong competition between overlapping watermarks. As one watermarker attempts to maintain its detection accuracy, the others’ detection accuracy decreases. SIR-SIR is not listed since the SIR does not allow the user to choose a key to create a different distribution, and therefore the watermark collision will not occur on SIR-SIR. However, SIR watermarks are still vulnerable to collision attacks using other watermark methods.